Mars Corporation uses the completed-contract method. At the end of the first year of a $9,000,000 contract, the following information is available:

Costs to date: $2,000,000
Estimated costs to complete 6,000,000
Progress billings during the year 1,800,000
Cash collected during the year 1,500,000


In the first year, Mars should recognize gross profit of
Answer #1

Answer:

First method:

Mars should recognize the gross profit, the method should be used completed contract method

Calculations are as:

Formula:

Gross profit= contact price-total estimated cost

Total estimate cost= cost to date+estimated cost to complete

Total estimated cost=2000,000+6000,000

=8000,000

Gross profit = 9000,000-800,000

=1000,000(completed contact method)

Second method:

​​​​​​Mars should recognize the gross profit, the method should be use percentage of completion method

Calculations are as:

Formula:

(Cost to date/Total estimated cost)*(contract price-total estimated cost)

Total estimate cost=cost to date+estimated cost to complete

Total estimated cost=2000,000+6000,000=8000,000

Gross profit=(2000,000/8000,000)*(9000,000-8000,000)

=0.25*1000,000

=250000( method used percentage completion method)